Creating equations:
the slope is 2 and the graph passes through ordered pair (1,-5).. what is the equation? and please write the explanition. Show your work.
the slope is 0 and the graph passes through ordered pair (0,-7).. what is the equation? and please write the explanition. Show your work.
the slope is undefined and the graph passes through ordered pair (12,0).. what is the equation? and please write the explanition. Show your work.
the slope is -1/3 and the graph passes through ordered pair (-6,-7).. what is the equation? and please write the explanition. Show your work.
the equation is parallel to the following equation y=-2/3x+5, and goes through the following ordered pairs (6,-7) what is the equation? and please write the explanition. Show your work.
the equation is perpendicular to the following equation y=-2/3x+5, and goes through the following ordered pairs (6,-7)

Answer by palanisamy(496)

1) the slope is 2 and the graph passes through ordered pair (1,-5)..
The equation of the line is
(y+5) = 2(x-1)
y+5 = 2x-2
y = 2x-7
2) the slope is 0 and the graph passes through ordered pair (0,-7)..
The equation of the line is
(y+7) = 0(x-0)
y+7=0
y=-7
3) the slope is undefined and the graph passes through ordered pair (12,0)..
The equation of the line is
x-12=0
x=12
4) the slope is -1/3 and the graph passes through ordered pair (-6,-7)..
The equation of the line is
(y+7) = -1/3(x+6)
y+7=-x/3-2
y=-x/3 -9
5) the equation is parallel to the following equation y=-2/3x+5, and goes through the following ordered pairs (6,-7)
Let the equation of the required line be
y = (-2/3)x+k ...(1)
The point (6,-7) lies on it
So, -7 = (-2/3)*6+k
-7 +4=k
k=-3
So, (1) => the equation of the required line be
y = (-2/3)x - 3
6)the equation is perpendicular to the following equation y=-2/3x+5, and goes through the following ordered pairs (6,-7)
The slope of the given line = -2/3
The slope of the perpendicular line is 3/2
The equation of the required line is
y+7 = (3/2)(x-6)
y = (3/2)x-9-7
y = (3/2)x-16